# Aligning Individual and Collective Objectives in Multi-Agent Cooperation

### v1

@@ Appendix B Additional Experiment Results
B.1 Result Visualization of Two-player Public Goods Game
@@
Appendix C Experiment Details
C.1 Social Dilemma Games: Cleanup and Harvest
SSD games (Vinitsky et al., 2019) implements the Harvest and Cleanup as grid world games.The agents use partially observed graphics observation, which contains a grid of $15\times 15$ centered on themselves. Compared to the original Cleanup and Harvest games, SSD games add a fire beam mechanic, where agents can fire on the grid to hit other partners. An agent will gain 1 reward if harvesting an apple and -1 reward if firing a beam. Besides, being hit by a beam will result in a 50 individual reward loss. Under this mechanic, selfish agents can easily use fire to prevent others from harvesting apples to gain more individual rewards but harm social welfare.
We utilized PPO algorithm in stable-baselines3 (Raffin et al., 2021) to implement the baselines and our methods, with all the agents using separated policy parameters for Simul-Ind and sharing same policy parameters for other experiments. For SVO, we modify the individual reward to be $r_{i}-\alpha(1-actan\left(\frac{sum_{j,j\neq i}r_{j}}{r_{i}}\right)$. To employ CGA and AgAs to PPO training process, we compute both individual and collective PPO-Clip policy losses and subsequently utilize them to calculate the adjusted gradient through automatic differentiation. We don’t make changes to the critic loss nor the critic net optimization process.
The hyper-parameters for PPO training are as follows.
@@ C.1 Social Dilemma Games: Cleanup and Harvest
C.2 Mixed-motive StarCraft II
SMAC, often employed as a testbed within the sphere of cooperative MARL, is characterized by the shared reward system among players. Within this paper, we present the MMM2 map transformed as a versatile testbed designed for mixed-motive cooperation. Specifically, MMM2 map encompasses 10 agents and 12 adversaries as enumerated below:
@@ C.2 Mixed-motive StarCraft II
Experiment settings. Simul-Ind leverages IPPO with recurrent policies and distinct policy networks. Conversely, Simul-Co employs MAPPO with recurrent policies and consolidated policy networks. The implementation of the IPPO and MAPPO algorithms presented in this paper is founded on the methodology detailed in (Yu et al., 2022). We utilize gradient adjustment optimization methods, such as CGA and AgAs, derived from MAPPO, implementing gradient adjustments as outlined in Algorithm 1.
The hyper-parameters for PPO-based training are as follows.

### v2

@@ Appendix D Additional Experiment Results
D.1 Result Visualization of Two-player Public Goods Game
@@
Appendix E Experiment Details
E.1 Sequential Social Dilemma Games: Cleanup and Harvest
SSD games [Vinitsky et al., 2019] implements the Harvest and Cleanup as grid world games.The agents use partially observed graphics observation, which contains a grid of $15\times 15$ centered on themselves. Compared to the original Cleanup and Harvest games, SSD games add a fire beam mechanic, where agents can fire on the grid to hit other partners. An agent will gain 1 reward if harvesting an apple and -1 reward if firing a beam. Besides, being hit by a beam will result in a 50 individual reward loss. Under this mechanic, selfish agents can easily use fire to prevent others from harvesting apples to gain more individual rewards but harm social welfare.
We utilized PPO algorithm in stable-baselines3 [Raffin et al., 2021] to implement the baselines and our methods, with all the agents using separated policy parameters for Simul-Ind and sharing same policy parameters for other experiments. For SVO, we modify the individual reward to be $r_{i}-\alpha(1-actan\left(\frac{sum_{j,j\neq i}r_{j}}{r_{i}}\right)$. To employ CGA and AgAs to PPO training process, we compute both individual and collective PPO-Clip policy losses and subsequently utilize them to calculate the adjusted gradient through automatic differentiation. We don’t make changes to the critic loss nor the critic net optimization process.
Most experiments were conducted on a node with a Tesla V100 GPU (32GB memory) and 40 CPU cores. The hyper-parameters for PPO training are as follows.
@@ E.1 Sequential Social Dilemma Games: Cleanup and Harvest
E.2 Selfish-MMM2
Figure 5: Semantic Diagram of MMM2 map in SMAC
SMAC, often employed as a testbed within the sphere of cooperative MARL, is characterized by the shared reward system among players. Within this paper, we present the MMM2 map transformed as a versatile testbed designed for mixed-motive cooperation. Specifically, as shown in Fig. 5, MMM2 map encompasses 10 agents and 12 adversaries as enumerated below:
@@ E.2 Selfish-MMM2
Experiment settings. Simul-Ind leverages IPPO with recurrent policies and distinct policy networks. Conversely, Simul-Co employs MAPPO with recurrent policies and consolidated policy networks. The implementation of the IPPO and MAPPO algorithms presented in this paper is founded on the methodology detailed in [Yu et al., 2022]. We utilize gradient adjustment optimization methods, such as CGA and AgAs, derived from MAPPO, implementing gradient adjustments as outlined in Algorithm 1.
Most experiments were conducted on a node with two NVIDIA GeForce RTX 3090 GPUs and 32 CPU cores. The hyper-parameters for PPO-based training are as follows.